Shift report 20020110 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Jan 10 2002 - 15:55:03 EST

  • Next message: Flemming Videbaek: "Fw: Controlled Access tomorrow"

    ------------------------------------------------------------
    Date: Thu Jan 10 15:55:03 2002, Shift: 08:00-16:00
    Supervisor: BB, with: YKL
    Log book#12 pages 40 - 41
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam for most of the shift(interurupted by polarization studies).
    
    2. Run 6339 Completed study of inelastic coincidences with L1 and R1 out of the coincidence)
       Run 6400 Completed data acquisition for MRS at 40 degrees and 4 kilogauss. FS was at 12 degrees and 1/4 B.
       Moved MRS to 45 degrees and started taking data(run 6401)
    
       Voted to dump beam when scaler 4 fell below 2k events/sec.
    
    3. No dtector problems.
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Jan 10 2002 - 15:55:34 EST